# A Review of Scalable and Privacy-Preserving Multi-Agent Frameworks for Distributed Energy Resource Control

### v1

@@ III-B3 Objective functions
$f_{\mathrm{env}}(\bm{x})=c_{s}p^{\mathrm{grid}}(\bm{x})+\sum_{i\in\mathcal{N}}\sum_{u\in\hat{\mathcal{N}}_{i}}g_{u,i}p_{u,i}^{\mathrm{fuel}}(\bm{x}_{u})$ (12)
where $p^{\mathrm{grid}}(\bm{x})$ denotes the total consumer power of grid electricity, multiplied by $c_{s}$ that denotes the carbon intensity of the grid electricity, $\hat{\mathcal{N}}_{i}$ denotes the set of agents connected to bus $i$, $p_{u,i}^{\mathrm{fuel}}(\bm{x}_{u})$ is the consumed fuels from other DER and non-DER sources, and $g_{u,i}$ denotes the carbon intensity of the specific fuel $u$ at bus $i$.
@@
$f_{\mathrm{curtail}}(\bm{x}^{\mathrm{pv}}_{\hat{\imath}})=||\bm{x}^{\mathrm{pv}}_{\hat{\imath}}-\bar{\bm{x}}^{\mathrm{pv}}_{\hat{\imath}}||_{2}^{2}+f^{\mathrm{pvg}}(s^{\mathrm{pv}}_{t})$ (14)
where $\bm{x}^{\mathrm{pv}}_{\hat{\imath}}\in\mathbb{R}^{T}$ and $\bar{\bm{x}}^{\mathrm{pv}}_{\hat{\imath}}\in\mathbb{R}^{T}$ denote the curtailed and original active power generations from the solar PV, respectively, and $f^{\textsc{pvg}}(s^{\mathrm{pv}}_{t})$ denotes the solar PV generation cost that can be described via a polynomial of the apparent power $s^{\mathrm{pv}}_{t}$, e.g., $f^{\mathrm{pvg}}(s^{\mathrm{pv}}_{t})=c_{1}^{\mathrm{pv}}(s^{\mathrm{pv}}_{t})^{2}+c_{2}^{\mathrm{pv}}s^{\mathrm{pv}}_{t}+c_{3}^{\mathrm{pv}}$, whose coefficients $c_{1}^{\mathrm{pv}}$, $c_{2}^{\mathrm{pv}}$, and $c_{3}^{\mathrm{pv}}$ can determined by curve fitting from the manufacturer. Broadly, curtailment cost objectives can be viewed as driving the states of grid-tied devices to desired values, similar to the requirement of returning a battery to its initial SoC at the end of the working period. Additionally, the competitive DER-level objectives also include the aggregated decision-making for a group of DERs, such as the bidding plans from distribution companies and DER aggregators [41] and the negotiation on locational marginal price from multiple prosumers [102].
III-C General Problem Formulation
@@
IV Scalable Methods
This section reviews state-of-the-art and fundamental scalable algorithms within multi-agent frameworks. We select representative works under each category of scalable multi-agent approaches and show their applications in DER control with highlighted key features (see Table I). At the end, we discuss related privacy leakage issues in these typical multi-agent frameworks for DER control.
IV-A Distributed and Decentralized Algorithms
TABLE I: Representative Work on Scalable Multi-Agent Frameworks and Their Applications in DER Control.
Method Reference Structure Applications Key Features
PGD [120] Smooth convex optimization 1-Prove convergence of gradient descent using nonconstant, long stepsize patterns, for smooth convex optimization; 2-via a computer-assisted analysis technique.
MARL [121] Decentralized (partially observable) Mobile power sources and repair crews 1-Formulate a resilience-driven dispatch problem; 2-a hierarchical MARL with embedded function encapsulating system dynamics.
MARL [124] Distributed Networked multi-agent problems 1-Cooperatively maximize the average of their entropy-regularized long-term rewards; 2-localized policy iteration algorithm that provably learns a near-globally-optimal policy using only local information.
           Not defined in the literature    AvgC: Average consensus    ADMM: Alternating direction method of multipliers    PGD: Projected gra-
           dient descent    MARL: Multi-agent reinforcement learning
IV-A1 Average consensus
Average consensus (AvgC) includes dynamic AvgC, where agents seek to compute the average of individual time-varying signals, and static AvgC, where agents reach the average of their initial values. The convergence of AvgC is first proved by DeGroot [125], then further studied by many researchers (see, e.g., [50, 112, 126]).
To provide a straightforward explanation, we refer to AvgC as the static one and introduce its theoretical foundations. AvgC-based algorithms are commonly used in multi-agent systems to collaboratively compute the average of agents’ local values. Suppose each agent has an initial scalar state $x_{i}^{0}$. The average consensus asymptotically converges to an “agreement,” e.g., a constant $c$, under suitable assumptions on the coefficients and graph connectivity. At the $\ell$th iteration, agent $\hat{\imath}$ updates its decision variable $x_{\hat{\imath}}^{\ell}\rightarrow x_{\hat{\imath}}^{\ell+1}$ by [125, 112]:
@@ IV-A1 Average consensus
where $\mathcal{B}_{\hat{\imath}}$ denotes the set of neighbors of agent $\hat{\imath}$, $\epsilon$ denotes the step size, and $\phi_{\hat{\imath}\hat{\jmath}}$ denotes the adjacency coefficient of the network, i.e., $\phi_{\hat{\imath}\hat{\jmath}}=0$ if $\hat{\jmath}\notin\mathcal{B}_{\hat{\imath}}$. Follow the distributed information exchange structure, the decision variables $x_{\hat{\imath}}^{\ell}$, $\forall\hat{\imath}\in\mathcal{B}$ converge to the averaged value $c=\sum_{\hat{\imath}=1}^{\hat{n}}x_{\hat{\imath}}^{0}/\hat{n}$, under balanced digraph and other numerical assumptions (see more details in [126, 50]). The scalable structure of distributed AvgC facilitates cooperative decision-making in networked multi-agent systems [127, 128, 129, 130, 131, 132, 133, 134, 135, 136], frequently used in networked control theory [128], communication networks [129, 130], load balancing [131, 132, 133, 134, 135], and formation control problems [136]. Variances of other AvgC-based algorithms include AvgC under asynchronous and time-varying environments [137, 138, 139], asymptotically accelerated AvgC using linear predictor [140], AvgC with quantization refinement to progressively reduce the quantization intervals during the algorithm convergence [141]. To summarize, the unique trait of AvgC ensures that all agents can reach an agreement based on their initial values/opinions, leading to well-suited deployment for DER control problems such as achieving optimal DER management for supply-demand balance [142, 143, 144, 133].
IV-A2 Alternating direction method of multipliers
@@
Since $f(\tilde{\bm{x}})$ and $g(\tilde{\bm{y}})$ have uncorrelated decision variables, the decomposability of ADMM allows $\tilde{\bm{x}}$ and $\tilde{\bm{y}}$ to be updated separately in a sequential (alternating) fashion. The distributed nature of ADMM enables parallel updates that are scalable for solving large-scale multi-agent optimization problems. Recently, ADMM has been extensively studied and improved with a number of generalizations, including approaches for tackling nonseparable optimization problem formulations [146, 147], nonconvex problems [148, 149, 150], as well as other heuristic ADMM-based approaches [151, 152]. ADMM-based methods are also being substantially studied for solving large-scale DER optimization problems, including the management of DERs with high uncertainty of power generation and load forecasts [116, 102], the decomposition of OPF with non-linear and non-convex formulations [153, 154], and asynchronous distributed optimization algorithms [155].
IV-A3 Projected gradient descent
@@
where $\alpha_{\hat{\imath}}$ and $\beta$ denote the primal and dual step sizes, respectively, $\mathcal{L}_{r}\left(\bm{x}_{1}^{\ell},\ldots,\bm{x}_{\hat{n}}^{\ell};\bm{\lambda}^{\ell}\right)$ denotes the relaxed Lagrangian function at the $\ell$th iteration, $\Pi_{\mathcal{X}_{\hat{\imath}}}[\cdot]$ denotes the Euclidean projection operator, and $\mathbb{R}^{+}$ denotes the positive real set. Regularization terms, such as the penalty $\frac{\rho}{2}\|\bm{D}\tilde{\bm{x}}+\bm{G}\tilde{\bm{y}}-\bm{h}\|_{2}^{2}$ in (17), can also be included in (19) to enforce the satisfaction of constraints and enhance convergence [52, 91].
PGD-based algorithms have been continuously improved on scalability, generality, and convergency for optimizing multi-agent systems. A series of recent findings include the regularized primal-dual subgradient method that can deal with non-separable objectives and constraints [52], shrunken primal-dual subgradient that eliminates the regularization errors [91], and shrunken primal-multi-dual subgradient that achieves two-facet scalable w.r.t. both the network dimension and the agent population size [156]. PGD-based (and gradient-based) approaches have been widely adopted for managing DER-populated power grids, examples include solving online load flow optimization problems [157], decentralized management of renewable generations and demand response [158], and voltage regulation with DERs [159].
IV-A4 Multi-agent reinforcement learning
@@
MARL involves the interactive decision-making of multiple agents that operate in a common environment [162, 163]. Many power system management problems, including DER control, can be cast into the realm of MARL, where various grid components, such as generators, controllers, or local operators, can act as independent agents and operate under the grid environment. In MARL, the $\hat{\imath}$th agent takes an action $a_{t}^{\hat{\imath}}\in\mathcal{A}_{\hat{\imath}}$, given the state $s_{t}$, and receives a reward $r_{t}^{\hat{\imath}}(s_{t},\{a_{t}^{\hat{\imath}}\}_{\hat{\imath}\in\mathcal{I}})$, then the system state $s_{t}$ transits into $s_{t+1}$. In power system applications, the states can include currents or voltages at different buses, real and reactive power demands, line flows, the status of DERs (e.g., battery energy level), transformer tap positions, etc. The actions can be taken on adjusting active/reactive power outputs, changing transformer tap settings, switching capacitors or reactors (on/off), adjusting load shedding levels, reconfiguring network topology, etc. Scalable MARL methods are powerful and promising tools for controlling DERs in large-scale power system networks with high-dimensional data streams [164, 165, 166, 167].
Based on the information exchange pattern between agents and the SO/coordinator, MARL algorithms can also be categorized into three representative types presented in Fig. 2. Specifically: 1) Centralized, also referred to as centralized training with decentralized execution. It assumes the existence of a central controller that can collect data from all agents, including their actions, rewards, and observations. Having a coherent view of the entire system, centralized settings significantly simplify the analysis [168, 169, 170]. However, centralized settings unavoidably suffer from scalability issues, such as the exponential growth of the joint action space. The scalability challenge motivates the development of decentralized or distributed structures that do not rely on a central controller; 2) Fully decentralized, where there is no direct exchange of information between any agents. Instead, each agent makes decisions independently based on its local observations, without any coordination and/or data aggregation. Fully decentralized structures largely enhance the scalability and eliminate peer-to-peer communications. However, they could suffer from non-convergence issues or delayed learning caused by the lack of a global view; 3) Distributed structure allows agents to communicate with others (e.g., neighbors) through a potentially time-varying communication network. Owing to the additional share of local information, distributed MARL structures benefit from better theoretical conciseness.

### v2

@@ 2.3.2 Constraints
DERs, such as solar PVs or ESSs, often provide energy or manage energy distribution to maintain comfort levels in buildings, particularly through HVACs. The operation of an HVAC system can also be interpreted as a type of thermostatically controlled load (TCL) that couples thermal dynamics (e.g., insulation, ambient temperature) and electricity consumption (e.g., On and Off cycles). For example, an HVAC can affect the indoor temperature according to the following linear dynamics [65]:
$\theta_{t}^{\mathrm{in}}=\theta_{t\!-\!1}^{\mathrm{in}}+\alpha_{h}(\theta_{t}^{\mathrm{out}}-\theta_{t\!-\!1}^{\mathrm{in}})+\beta_{h}x_{t}^{\mathrm{H}}$ (7)
where $x_{t}^{\mathrm{H}}$ denotes the active power consumption of the HVAC at time $t$, $\theta_{t}^{\mathrm{in}}$ and $\theta_{t}^{\mathrm{out}}$ denote the indoor and outdoor temperatures at time $t$, respectively. The parameters $\alpha_{h}$ and $\beta_{h}$ define the heat transfer properties of the environment and the thermal efficiency of the TCL appliance, respectively. The comfort zone constraint defines that the average temperature of the building should remain within the range of:
$\underline{\theta}\leq{\bm{\theta}}^{\mathrm{in}}\leq\overline{\theta}$ (8)
where ${\bm{\theta}}^{\mathrm{in}}\in\mathbb{R}^{T}$ denotes the indoor room temperature across $T$ time intervals, $\underline{\theta}$ and $\overline{\theta}$ denote the lower and upper bounds of the comfort zone, respectively.
@@ 2.3.3 Objective functions
$f_{\mathrm{env}}(\bm{x})=c_{s}p^{\mathrm{grid}}(\bm{x})+\sum_{i\in\mathcal{N}}\sum_{u\in\hat{\mathcal{N}}_{i}}g_{i,u}p_{i,u}^{\mathrm{fuel}}(\bm{x}_{u})$ (14)
where $p^{\mathrm{grid}}(\bm{x})$ denotes the total consumer power of grid electricity, multiplied by $c_{s}$ that denotes the carbon intensity of the grid electricity, $\hat{\mathcal{N}}_{i}$ denotes the set of agents connected to bus $i$, $p_{i,u}^{\mathrm{fuel}}(\bm{x}_{u})$ is the consumed fuels from other DER and non-DER sources, $\bm{x}_{u}$ denotes the decision variables of the $u$th fuel source, and $g_{i,u}$ denotes the carbon intensity of the specific fuel $u$ at bus $i$.
@@
The comfort to users based on electrical usage of common appliances, such as washing machines and HVAC systems, is also commonly considered a DER-level objective. The On-Off status of an HVAC can be controlled by end users to maintain room temperature within a specific comfort zone, i.e., $[\underline{\theta},\overline{\theta}]$. To this end, the indoor room temperature comfort objective can be enforced by:
$f_{\mathrm{HVAC}}(\bm{x}_{\hat{\imath}}^{\mathrm{H}})=\|\bm{\theta}^{\mathrm{in}}-\hat{\bm{\theta}}^{\mathrm{in}}\|_{2}^{2}$ (16)
where $\hat{\bm{\theta}}^{\mathrm{in}}\in\mathbb{R}^{T}$ denotes the desired room temperature across $T$ time intervals, e.g., the averaged temperature of $(\underline{\theta}+\overline{\theta})/2$.
Washing appliances, like dishwashers and clothes washers, provide comfort to users when tasks are completed by a specific time [65]. In other words, the cumulative power consumption of such machines must reach a certain energy threshold by the deadline, defined by:
$f_{\mathrm{machine}}(\bm{x}_{\hat{\imath}}^{\mathrm{M}})=|\sum_{t=1}^{T}(\bm{x}_{\hat{\imath},t}^{\mathrm{M}}\cdot\Delta t)-\hat{e}^{\mathrm{M}}_{T}|^{2}$ (17)
where $\bm{x}_{\hat{\imath}}^{\mathrm{M}}\in\mathbb{R}^{T}$ denotes the active power consumption of the $\hat{\imath}$th washing machine across $T$ time intervals, $\Delta t$ denotes the time interval length, and $\hat{e}^{\mathrm{M}}_{T}$ denotes the desired energy level at the end of the period.
@@
$f_{\mathrm{curtail}}(\bm{x}^{\mathrm{pv}}_{\hat{\imath}})=||\bm{x}^{\mathrm{pv}}_{\hat{\imath}}-\bar{\bm{x}}^{\mathrm{pv}}_{\hat{\imath}}||_{2}^{2}+f^{\mathrm{pvg}}(\bm{s}^{\mathrm{pv}})$ (18)
where $\bm{x}^{\mathrm{pv}}_{\hat{\imath}}\in\mathbb{R}^{T}$ and $\bar{\bm{x}}^{\mathrm{pv}}_{\hat{\imath}}\in\mathbb{R}^{T}$ denote the curtailed and original active power generation from the solar PV, respectively, and $f^{\textsc{pvg}}$ ($\bm{s}^{\mathrm{pv}}$) denotes the solar PV generation cost that can be described via a polynomial of the apparent power $\bm{s}^{\mathrm{pv}}\in\mathbb{R}^{T}$, e.g., $f^{\mathrm{pvg}}(s^{\mathrm{pv}}_{t})=c_{1}^{\mathrm{pv}}(s^{\mathrm{pv}}_{t})^{2}+c_{2}^{\mathrm{pv}}s^{\mathrm{pv}}_{t}+c_{3}^{\mathrm{pv}}$, whose coefficients $c_{1}^{\mathrm{pv}}$, $c_{2}^{\mathrm{pv}}$, and $c_{3}^{\mathrm{pv}}$ can determined by curve fitting from the manufacturer. Broadly, curtailment cost objectives can be viewed as driving the states of grid-tied devices to desired values, similar to the requirement of returning a battery to its initial SoC at the end of the working period. Additionally, the competitive DER-level objectives also include the aggregated decision-making for a group of DERs, such as the bidding plans from distribution companies and DER aggregators [56] and the negotiation on locational marginal price from multiple prosumers [112].
2.4 General Problem Formulation
@@
3 Scalable Methods
This section reviews state-of-the-art and fundamental scalable algorithms within multi-agent frameworks. We select representative works under each category of scalable multi-agent approaches and show their applications in DER control with highlighted key features (see Table LABEL:table_scalable_alg). At the end, we discuss related privacy leakage issues in these typical multi-agent frameworks for DER control.
3.1 Distributed and Decentralized Algorithms
3.1.1 Average consensus
Average consensus (AvgC) includes dynamic AvgC, where agents seek to compute the average of individual time-varying signals, and static AvgC, where agents reach the average of their initial values. The convergence of AvgC is first proved by DeGroot [122], then further studied by many researchers (see, e.g., [68, 123, 124]). To provide a straightforward explanation, we refer to AvgC as the static one and introduce its theoretical foundations. AvgC-based algorithms are commonly used in multi-agent systems to collaboratively compute the average of agents’ local values. Suppose each agent has an initial scalar state $x_{i}^{0}$. The average consensus asymptotically converges to an “agreement,” e.g., a constant $c$, under suitable assumptions on the coefficients and graph connectivity. At the $\ell$th iteration, agent $\hat{\imath}$ updates its decision variable $x_{\hat{\imath}}^{\ell}\rightarrow x_{\hat{\imath}}^{\ell+1}$ by [122, 123]:
@@ 3.1.1 Average consensus
where $\mathcal{B}_{\hat{\imath}}$ denotes the set of neighbors of agent $\hat{\imath}$, $\epsilon$ denotes the step size, and $\phi_{\hat{\imath}\hat{\jmath}}$ denotes the adjacency coefficient of the network, i.e., $\phi_{\hat{\imath}\hat{\jmath}}=0$ if $\hat{\jmath}\notin\mathcal{B}_{\hat{\imath}}$. Follow the distributed information exchange structure, the decision variables $x_{\hat{\imath}}^{\ell}$, $\forall\hat{\imath}\in\mathcal{B}$ converge to the averaged value $c=\sum_{\hat{\imath}=1}^{\hat{n}}x_{\hat{\imath}}^{0}/\hat{n}$, under balanced digraph and other numerical assumptions (see more details in [124, 68]). The scalable architecture of distributed AvgC enables cooperative multi-agent decision-making [125, 126, 127, 128, 129, 130, 131, 132, 133]. These AvgC methods are widely utilized in diverse fields that pertinent to multi-agent-based DER control problems, including networked control theory [134], communication networks [126, 127]. For smart grid applications, AvgC-based methods have gained prominence, especially in the control of grid-tied energy resources [128, 129, 130, 131, 132, 135]. The unique trait of AvgC that ensures that all agents can reach an agreement based on their initial values/opinions offers the opportunity for well-suited deployment of DER control problems, such as achieving optimal DER management for supply-demand balance [136, 137, 138, 130]. Therefore, AvgC is efficient for distributed coordination and primarily used for tasks like distributed averaging, decentralized estimation, and synchronization in networked multi-agent systems. However, AvgC-based methods are often limited to scenarios where all agents must reach the same consensus value. Additionally, future research is needed to accelerate convergence in large networks or those with weak connectivity and to reduce sensitivity to time delays and changes in network topology. Variations of AvgC-based algorithms have been developed to handle asynchronous and time-varying environments [139, 140, 141], accelerate convergence through linear predictors [142], and incorporate quantization techniques that refine intervals as the algorithm progresses [143]. These new techniques are crucial for the research and development in the power and energy field, as power systems inherently exist as networked systems in dynamic environments.
3.1.2 Alternating direction method of multipliers
@@
Since $f(\tilde{\bm{x}})$ and $g(\tilde{\bm{y}})$ have uncorrelated decision variables, the decomposability of ADMM allows $\tilde{\bm{x}}$ and $\tilde{\bm{y}}$ to be updated separately in a sequential (alternating) fashion. The distributed nature of ADMM enables scalability in solving large-scale multi-agent optimization problems. By decomposing large problems into smaller sub-problems, ADMM-based approaches allow individual agents to solve these sub-problems locally. They are suitable for distributed optimization problems such as distributed machine learning [145], large-scale networked systems [146], and resource allocation problems in power grids and communication networks [112]. Despite the merits, ADMM-based methods can suffer from high communication overhead between agents due to the need for iterative message exchanges. Recently, ADMM has been extensively studied and improved with a number of generalizations, including approaches for tackling nonseparable optimization problem formulations [147, 148], nonconvex problems [149, 150, 151], as well as other heuristic ADMM-based approaches [152, 153]. Therefore, ADMM-based methods are well-suited for solving large-scale DER optimization problems, including the management of DERs with high uncertainty of power generation and load forecasts [154, 112], the decomposition of OPF with non-linear and non-convex formulations [155, 156], and asynchronous distributed optimization algorithms [157].
3.1.3 Projected gradient descent
@@
where $\alpha_{\hat{\imath}}$ and $\beta$ denote the primal and dual step sizes, respectively, $\mathcal{L}_{r}(\bm{x}_{1}^{\ell},\ldots$, $\bm{x}_{\hat{n}}^{\ell};\bm{\lambda}^{\ell})$ denotes the relaxed Lagrangian function at the $\ell$th iteration, $\Pi_{\mathcal{X}_{\hat{\imath}}}[\cdot]$ denotes the Euclidean projection operator, and $\mathbb{R}^{+}$ denotes the positive real set. Regularization terms, such as the penalty $\frac{\rho}{2}\|\bm{D}\tilde{\bm{x}}+\bm{G}\tilde{\bm{y}}-\bm{h}\|_{2}^{2}$ in (21), can also be included in (23) to enforce the satisfaction of constraints and enhance convergence [70, 101].
PGD-based methods are straightforward to implement and computationally efficient for large-scale multi-agent problems. The PGD allows agents to handle local constraints individually, making it a good fit for decentralized settings. However, PGD can suffer from slow convergence when addressing non-convex problems or poorly conditioned constraints, as it is sensitive to step size choices and requires careful tuning for optimal performance. Additionally, agents must perform a projection step, which can be computationally expensive for certain constraint sets. To enhance scalability, generality, and convergency in optimizing multi-agent systems, PGD-based algorithms have been continuously improved to include a series of recent findings, such as the regularized primal-dual subgradient method that can deal with non-separable objectives and constraints [70], shrunken primal-dual subgradient that eliminates the regularization errors [101], and shrunken primal-multi-dual subgradient that achieves two-facet scalable w.r.t. both the network dimension and the agent population size [66]. PGD-based (and gradient-based) approaches have been widely adopted for managing DER-rich power grids, examples include solving online load flow optimization problems [158], decentralized management of renewable generation and demand response [159], and voltage regulation with DERs [160].
3.1.4 Multi-agent reinforcement learning
@@
MARL involves the interactive decision-making of multiple agents that operate in a common environment [163, 164]. Many power system management problems, including DER control, can be cast into the realm of MARL, where various grid components, such as generators, controllers, or local operators, can act as independent agents and operate under the grid environment. In MARL, the $\hat{\imath}$th agent takes an action $a_{t}^{\hat{\imath}}\in\mathcal{A}_{\hat{\imath}}$, given the state $s_{t}$, and receives a reward $r_{t}^{\hat{\imath}}(s_{t},\{a_{t}^{\hat{\imath}}\}_{\hat{\imath}\in\mathcal{I}})$, then the system state $s_{t}$ transits into $s_{t+1}$. In power system applications, the states can include currents or voltages at different buses, real and reactive power demands, line flows, the status of DERs (e.g., battery energy level), transformer tap positions, etc. The actions can be taken on adjusting active/reactive power outputs, changing transformer tap settings, switching capacitors or reactors (on/off), adjusting load shedding levels, reconfiguring network topology, etc.
MARL-based approaches are compatible with dynamic, uncertain, and complex environments where agents learn to make decisions through interactions with the environment. When predefined strategies fail in uncertain environments, MARL enables agents to adapt by learning from their actions and their consequences over time, capable of solving a wide range of dynamic tasks, including coordination, cooperation, and competition between agents. Based on the information exchange pattern between agents and the SO or coordinator, MARL algorithms can also be categorized into three representative types as presented in Fig. 2, including centralized (also referred to as centralized training with decentralized execution), decentralized, and distributed structures (also referred to as decentralized setting with networked agents), see [163] for more details. Scalable MARL methods are powerful and promising tools for controlling DERs in large-scale power system networks with high-dimensional data streams [165, 166, 167, 168]. However, MARL-based methods commonly face high computational complexity and slow convergence in large-scale systems due to the curse of dimensionality and exploration-exploitation trade-offs. Besides, coordination between agents is challenging in dynamic power systems, especially in partially observable environments. Another future direction involves addressing the requirement for extensive training data, as learned policies may not generalize well to new scenarios, such as cyber and physical attack situations in power systems [169].
Remark 1
Table 1: Representative Work on Scalable Multi-Agent Frameworks and Their Applications in DER Control.
Method Ref. Structure Applications Key Features
PGD [175] Smooth convex optimization 1-Prove convergence of gradient descent using nonconstant, long stepsize patterns, for smooth convex optimization; 2-computer-assisted analysis.
MARL [176] Hierarchical (partially observable) Mobile power sources and repair crews 1-Formulate a resilience-driven dispatch problem; 2-a hierarchical MARL with embedded function encapsulating system dynamics.
MARL [179] Distributed Networked multi-agent problems 1-Cooperatively maximize the average of their entropy-regularized long-term rewards; 2-localized policy iteration algorithm that provably learns a near-globally-optimal policy using only local information. 11footnotemark: 1
Table 1: (Continued)
Not defined in the literature    Ref.: Reference    AvgC: Average consensus    ADMM: Alternating direction method of multipliers    PGD: Projected gradient descent    MARL: Multi-agent reinforcement learning
3.2 Privacy Leakages
To ensure the warranted use of private information from all stakeholders, it is crucial to synthesize privacy preservation techniques into the design of scalable DER control strategies. Toward this goal, we summarize the typical adversaries/threats in multi-agent computing frameworks, including external eavesdroppers, honest-but-curious agents, and the SO and/or coordinators/aggregators, each representing a distinct type of threat with different attack vectors. By examining these three adversaries, we cover a broad spectrum of external, internal, and hierarchical privacy risks, offering a comprehensive understanding of the threats in multi-agent systems. This helps guide the design of privacy-preserving multi-agent frameworks that require privacy guarantees in different operational scenarios.
3.2.1 External eavesdroppers
3.2.2 Honest-but-curious agents
3.2.3 The system operator and/or coordinators/aggregators